QUESTION 1 Find the equation of the regression line. x | 52 60 54 53 5556 y1308 338 308 298 298 308 aragraph M Arial M 3 (12pt) ?I :-- E-Ti, Path Click Save and Submit to save and submit Click Save A Anstwers to save all answers
Answer #1

Solution

Back-up Theory

The linear regression model Y = ?0 + ?1X + ?, ………………………………………..(1)

where ? is the error term, which is assumed to be Normally distributed with mean 0 and variance ?2.

Estimated Regression of Y on X is given by: Y = ?0cap + ?1capX, ………………………….(2)

where

?1cap = Sxy/Sxx and ?0cap = Ybar – ?1cap.Xbar..……………………………………………..(3)

Mean X = Xbar = (1/n)sum of xi ………………………………………….……………….(4)

Mean Y = Ybar = (1/n)sum of yi ………………………………………….……………….(5)

Sxx = sum of (xi – Xbar)2 …………………………………………………..………………………………..(6)

Syy = sum of (yi – Ybar)2 …………………………………………………..………………………………..(7)

Sxy = sum of {(xi – Xbar)(yi – Ybar)} …………………………………………………………………….………(8)

All above sums are over i = 1, 2, …., n

n = sample size ………………………………………………………………………………(9)

Calculations:

Given data

i

xi

yi

1

52

308

2

60

338

3

54

308

4

53

298

5

55

298

6

56

308

Summary of Excel Calculations

n

6

xbar

55

ybar

309.6667

Sxx

40

Syy

1083.3333

Sxy

170

?1cap

4.25

?0cap

75.9167

Thus, the regression equation is: y = 75.92 + 4.25x ANSWER

DONE